Two sides of a triangle have the same length. The third side is 20 cm longer than each of the other side. The perimeter is no less than 95 cm. At least how long are the sides of the triangle?

Answer by ikleyn(52775)

.

The sides are x, x and (x+20) centimeters.


The perimeter inequality is


    x + x + (x+20) >= 95.


It implies


    3x >= 95 - 20 = 75,

or

     x >= 75/3 = 25.


ANSWER.  Two shorter sides of same length are at least 25 cm long.

         The third side is at least 25 cm long.

Solved, answered and explained.